OpenStudy (anonymous):

Solve, |x + 1/3|=|x-3|

OpenStudy (anonymous):

I don't think there is a solution for this, lol

OpenStudy (anonymous):

If you graph both sides of the equation as separate functions on the same graph, they never intersect because they have the same slope.

OpenStudy (anonymous):

i agree with ambius

OpenStudy (anonymous):

Hmm okay, thank you. My book claims that the answer is 4/3 but I got no solution myself. Thank you

OpenStudy (anonymous):

oh wait,

OpenStudy (anonymous):

did yo try to square both side first?

OpenStudy (anonymous):

No thats the problem as it was written, I do nothing to it.

OpenStudy (anonymous):

\[x+\frac{ 1 }{ 3 }=-\left| x-3 \right|\] \[-x-\frac{1}{3}=\left|x-3\right|\] \[-x-\frac{1}{3}=x-3\] \[-2x=-\frac{8}{3}\] \[x=\frac{8}{6}\] \[x=\frac{4}{3}\]

OpenStudy (anonymous):

That's called doing maths.

OpenStudy (anonymous):

Dat it is, thank you good sir or madam.

OpenStudy (anonymous):

then u get: \[x^2+\frac{2x}{3}+\frac{1}{9}=x^2-6x+9\]

OpenStudy (anonymous):

that's lead to \(x=4/3\)

OpenStudy (anonymous):

way to go! I just got schooled in absolutes, lol